2 card are drawn from a pack of well shuffled 52 cards one by one with replacement .getting an ace or spade is considered success.find probability distribution for number of success

Grade:12

1 Answers

Yash Kawdiya
6 years ago
No. of Ace in cards = 4No. of Spade in cards = 13Total No. of required Cards = 17But there is a Ace of Spade which is repeated so we have to subtract it.Therefore Total No. of required Cards = 17 -1 = 16Total no. Of cards = 52Probablity = 16/52 = 0.3076But the Probablity is of 2 cards so it becomes double.Therefore final probablity = 0.3076 × 2 = 0.615Final ans = 0.615